# This work was supported by the Center for Identification Technology Research and the National Science Foundation, under Grants 2413228 and 2501916

### v1

@@ VI-A Spatial
From background research, we expected to see that a person‘s gaze would be less concentrated on their partner’s face when speaking and more so when they were listening [7], [8]. This is important as our expectation being wrong could indicate that our PoG estimation method is not sufficiently accurate.
@@ VII Detector
When designing the detector, we had two main considerations: the real-time requirement and our relatively limited amount of data. Given those requirements and based on testing, we decided to go with a simple 2D CNN architecture, which performed just as well or better than the other architectures we tested while remaining quick and computationally cheap. In the spirit of being thorough, we tested a transformer model, which we expected to do poorly given our small amount of data. To our surprise, it performed on par with the 2D CNN, but it was not nearly as fast or computationally cheap. It is worth noting that the use of a 2D CNN for time-series data is slightly unconventional, but we believe that the way in which we structured the input data allowed us to get good results.
@@ VII-A Features
Our system uses a total of 298 features, all of which, with the exception of the speaking context information, are derived from the six carefully chosen facial landmarks. For each of these six facial landmarks, the unit vector and magnitude component are calculated from the PoG to it; additionally, from the magnitude component of each facial landmark, a spectrogram is generated. The speaking context information is a crucial feature that allows the model to go beyond just finding mistakes in the gaze. With its inclusion the model can relate the PoG to who is speaking in the conversation allowing the model to gain access to the deeper patterns within social interaction that deepfakes are unable to mimic, e.g., gaze being used to precipitate turn taking during conversation [9] or how a speakers gaze wanders off their partners face and a listeners gaze is more attentive [7], [8].
@@
The locations and MediaPipe indices of the six facial landmarks can be seen in Fig. 5. The landmarks were chosen to capture the facial regions used in Vehlen et al. [8] while using the least amount of points. We decided to go with regions as defined in Vehlen et al. [8] facial regions because its 4 regions are well defined and capture meaningful data. Vehlen et al. [8] finds that roughly 70% of the gaze is concentrated on the eyes, mouth, and nose regions, while 10 to 15% falls on the rest of the face, with the remaining 10 to 15% missing or not falling on the face. Additionally due to the average 42 millimeters of error in the PoG estimation it would be unable to make use of more complicated facial regions.
@@ VII-B Hyperparameters
With these two considerations in mind, we determined an optimal sequence length of 1,800, which is about 1 minute at the average fps of 29. We determined that this was the optimal sequence length for a number of reasons. It was the longest it could be before becoming incompatible with a real-time system. Increasingly longer sequences were significantly more computationally expensive while only offering marginal accuracy improvements. This length is significantly more accurate than shorter lengths while only modestly increasing computational expense. Lastly, at this length, the sequences can be overlapped with their neighbors and used for voting, improving accuracy with little to no cost. This brings us to the stride setting, which controls the amount of overlap between sequences, and while not strictly a hyperparameter, it affects the model in a very similar way.
@@ VII-D Results
As can be seen in Fig. 8, our model is able to generalize well across all training validation splits, with the most accurate split having an accuracy of 88% and the least accurate split having an accuracy of 72%, an average validation accuracy of 82% when trained on both deepfake methods, and an average accuracy of 99% and 73% when using just FaceShifter and DeepFaceLive, respectively.
@@ VII-D1 Voting
TABLE III: The increase in accuracy provided by different amounts of voters.
@@
Table III shows the effect that voting has on the model’s accuracy with and without including FaceShifter in the dataset, which was done because it has a 99% accuracy. As can be seen in the table, the addition of voting does not have a large effect, and there is no noticeable difference between soft and hard. When it comes to an actual implementation, any voter amount over 10 would be inefficient, as each voter adds an additional 6 seconds of wait time before a prediction can be made.
@@ VIII Conclusion
Despite the success of the pipeline, there are a few limitations that should be mentioned. Although the PoG estimation system had no problem with glasses, if even the slightest amount of glare was present, it would give wildly inaccurate estimates. We compensated for this by using external lights and ensuring that the screen brightness was set correctly, but outside of a controlled environment, this would likely be impossible. Additionally, to our knowledge, there is no gaze tracking software capable of accurate estimation while the eyes are occluded. A possible partial remedy to this could be a system that would detect said occlusions/glare and prompt the user to correct it. Another limitation that was a result of time constraints was that we were unable to ensure that individuals with strabismus or social differences, which can cause differences in gaze [45], were properly represented in our dataset, meaning that these individuals could be falsely classified as deepfakes. This problem could be solved by the collection of a larger dataset where a portion of the subjects were sought out based on having the aforementioned characteristics. In addition to concerns that could be remedied, there are several avenues for future work. Such as exploring the creation of a system that could work in video call environments with more than two people, or the inclusion of additional biometrics, such as facial expression, and the content of conversation, which may contain additional patterns that can be exploited for deepfake detection.

### v2

@@ VI-A Spatial
From background research, we expected to see that a person‘s gaze would be less concentrated on their partner’s face when speaking and more so when they were listening [dualeyetrackingconvo], [eyetrackingsetupevel]. This is important as our expectation being wrong could indicate that our PoG estimation method is not sufficiently accurate.
@@ VII Detector
When designing the detector, we had two main considerations: the real-time requirement and our relatively limited amount of data. Given those requirements and based on testing, we decided to go with a simple 2D CNN architecture which uses a streaming inference mechanism to get input in near real-time. Our 2D CNN performed just as well or better than the other architectures we tested while remaining quick and computationally cheap. In the spirit of being thorough, we tested a transformer model, which we expected to do poorly given our small amount of data. To our surprise, it performed on par with the 2D CNN, but it was not nearly as fast or computationally cheap. It is worth noting that the use of a 2D CNN for time-series data is slightly unconventional, but its underlying architecture fits our needs. Its architecture allows the model to automatically relate all six of our feature streams together with a single light weight model through the use of 2D convolutions. As opposed to temporal CNNs whose 1D convolutions would necessitate additional steps in order to relate the six feature streams to one and other.
@@ VII-A Features
Our system uses a total of 298 features, all of which, with the exception of the speaking context information, are derived from the six carefully chosen facial landmarks. For each of these six facial landmarks, the unit vector and magnitude component are calculated from the PoG to it; additionally, from the magnitude component of each facial landmark, a spectrogram is generated. The speaking context information is a crucial feature that allows the model to go beyond just finding mistakes in the gaze. With its inclusion the model can relate the PoG to who is speaking in the conversation allowing the model to gain access to the deeper patterns within social interaction that deepfakes are unable to mimic, e.g., gaze being used to precipitate turn taking during conversation [gazesignals] or how a speakers gaze wanders off their partners face and a listeners gaze is more attentive [dualeyetrackingconvo], [eyetrackingsetupevel].
@@
The locations and MediaPipe indices of the six facial landmarks can be seen in Fig. 5. The landmarks were chosen to capture the facial regions used in Vehlen et al. [eyetrackingsetupevel] while using the least amount of points. We decided to go with regions as defined in Vehlen et al. [eyetrackingsetupevel] facial regions because its 4 regions are well defined and capture meaningful data. Vehlen et al. [eyetrackingsetupevel] finds that roughly 70% of the gaze is concentrated on the eyes, mouth, and nose regions, while 10 to 15% falls on the rest of the face, with the remaining 10 to 15% missing or not falling on the face. Additionally due to the average 42 millimeters of error in the PoG estimation it would be unable to make use of more complicated facial regions.
@@ VII-B Hyperparameters
With these two considerations in mind, we determined an optimal sequence length of 1,800, which is about 1 minute at the average fps of 29. Through testing we determined that this was the optimal sequence length for a number of reasons. It was the longest it could be before becoming incompatible with a real-time system. Increasingly longer sequences were significantly more computationally expensive while only offering marginal accuracy improvements. This length is significantly more accurate than shorter lengths while only modestly increasing computational expense. Lastly, at this length, the sequences can be overlapped with their neighbors and used for voting, improving accuracy with little to no cost. This brings us to the stride setting, which controls the amount of overlap between sequences, and while not strictly a hyperparameter, it affects the model in a very similar way.
@@ VII-D Results
TABLE III: Metrics at different level of gaussian noise perturbations.
Level of perturbation Metric
AUC EER Accuracy
None 88.20% 17.95% 82.52%
1 cm 89.12% 19.95% 83.64%
2 cm 85.03% 22.73% 81.02%
5 cm 83.69% 25.84% 79.44%
As can be seen in Fig. 8, our model is able to generalize well across all training validation splits, with the most accurate split having an accuracy of 88% and the least accurate split having an accuracy of 72%, an average validation accuracy of 82% when trained on both deepfake methods, and an average accuracy of 99% and 73% when using just FaceShifter and DeepFaceLive, respectively. Finally in order to better understand the 4.2 cm of error in the PoG calculation we perturbed the PoG using gaussian noise at three different levels to replicates an additional 1 cm, 2 cm, and 5 cm of error in the PoG estimation. The results of this can be seen in in table  III. As can be seen these perturbations did not have a large effect e.g. at 5 cm of perturbation the area under curve dropped by less than 5%. Given how small these reduction are we believe that the 4.2 cm of error in the PoG tracker does not excessively reduce our models accuracy.
@@ VII-D1 Voting
TABLE IV: The increase in accuracy provided by different amounts of voters.
@@
Table IV shows the effect that voting has on the model’s accuracy with and without including FaceShifter in the dataset, which was done because it has a 99% accuracy. As can be seen in the table, the addition of voting does not have a large effect, and there is no noticeable difference between soft and hard. When it comes to an actual implementation, any voter amount over 10 would be inefficient, as each voter adds an additional 6 seconds of wait time before a prediction can be made.
@@ VIII Conclusion
IX Future Work
Despite the success of the pipeline, there are a few limitations that would need to be addressed by future work. Although the PoG estimation system had no problem with glasses, if even the slightest amount of glare was present, it would give wildly inaccurate estimates. We compensated for this by using external lights and ensuring that the screen brightness was set correctly, but outside of a controlled environment, this would likely be impossible. But, to our knowledge, there is no gaze tracking software capable of accurate estimation while the eyes are occluded. A possible partial remedy to this could be a system that would detect said occlusions/glare and prompt the user to correct it. Additionally our current dataset has a few limitations. As a result of time constraints we were unable to ensure that individuals with strabismus or social differences, which can cause differences in gaze [eyemovementautistictraits], were properly represented in our dataset, meaning that these individuals could be falsely classified as deepfakes. Also as previously mentioned the dataset has a gender imbalance and finally the dataset was collected in a controlled lab environment meaning that diverse lighting and background conditions are not present. Future work that goes beyond the creation of a proof of concept pipeline will need to address these limitation, by conducting more extensive data collection to ensure the dataset is represents realistic lighting conditions and population demographics.
In addition to concerns that could be remedied by future work there are several ways to expand on this research. Such as including additional biometric like facial expression, and the content of conversation, which may contain additional patterns that can be exploited for deepfake detection. Another avenue for future work would be expanding the system to work in environments with more than two participants, this would be very valuable because the majority of calls include more than two participants. This new system would offer unique challenges and opportunities not present in the current system. While it would be easy to determine where the listeners are looking as the speaker would take up their entire screen, a feature already common in video conference apps. We believe that it would be difficult given the technical limitations of real-time point of gaze detection to discern the subtleties of gaze when the speaker is looking at multiple people who have been shrunk to fit on screen at once. On the other hand the fact that there are more than two participates could allow for their comparison to find outliers, which would give us another method of detection possibly improving our results. Lastly future work could try to incorporate non biometric detection methods such as water marking [xinliao2], [xinliao3], [xinliao4] or signal noise based detection [xinliao].